# Bozen-Bolzano, IT SkillFormer

### v2

@@ Abstract
Assessing human skill levels in complex activities is a challenging problem with applications in sports, rehabilitation, and training. In this work, we present SkillFormer, a parameter-efficient architecture for unified multi-view proficiency estimation from egocentric and exocentric videos. Building on the TimeSformer backbone, SkillFormer introduces a CrossViewFusion module that fuses view-specific features using multi-head cross-attention, learnable gating, and adaptive self-calibration. We leverage Low-Rank Adaptation to fine-tune only a small subset of parameters, significantly reducing training costs. In fact, when evaluated on the EgoExo4D dataset, SkillFormer achieves state-of-the-art accuracy in multi-view settings while demonstrating remarkable computational efficiency, using 4.5x fewer parameters and requiring 3.75x fewer training epochs than prior baselines. It excels in multiple structured tasks, confirming the value of multi-view integration for fine-grained skill assessment.
@@ 1 Introduction
To address this, we propose SkillFormer, a lightweight and efficient architecture for multi-view skill assessment. SkillFormer processes synchronized egocentric and exocentric video streams using a TimeSformer backbone, enhanced by a novel CrossViewFusion module that leverages cross-attention and learnable gating. To enable efficient fine-tuning, it incorporates Low-Rank Adaptation (LoRA). This design achieves competitive overall accuracy while substantially improving computational efficiency through reduced trainable parameters and faster training convergence.
@@ 2 Background and Related Works
Action Quality Assessment (AQA) and Proficiency Estimation focus on evaluating the quality of human actions, with applications across sports, rehabilitation, and skill training [Zhou et al.(2024)Zhou, Cai, Wang, Shum, and Liang]. Traditional approaches relied on single-modality inputs, using handcrafted features or 3D CNNs like C3D and I3D to extract spatiotemporal representations [Parmar and Morris(2019), Carreira and Zisserman(2017), Tran et al.(2015)Tran, Bourdev, Fergus, Torresani, and Paluri]. More recent methods adopt transformer-based models such as TimeSformer, Video Swin, and ActionFormer to capture long-range dependencies and fine-grained motion dynamics [Bertasius et al.(2021)Bertasius, Wang, and Torresani, Liu et al.(2021)Liu, Ning, Cao, Wei, Zhang, Lin, and Hu, Zhang et al.(2022)Zhang, Wu, and Li]. Specialized architectures like GDLT decouple grade-aware features via decoder attention [Xu et al.(2022)Xu, Zeng, and Zheng], while CoRe introduces contrastive regression for more precise skill scoring [Yu et al.(2021)Yu, Rao, Zhao, Lu, and Zhou]. Nonetheless, skill classification remains underexplored, particularly in multi-view settings.
Multi-modal AQA improves robustness by integrating complementary signals—RGB, flow, audio, depth, or pose. Skating-Mixer [Xia et al.(2022)Xia, Zhuge, Geng, Fan, Wei, He, and Zheng] and PAMFN [Zeng and Zheng(2024)] incorporate audio-visual cues via staged fusion. Others combine RGB and depth for industrial tasks [Bianchi and Lanz(2024)], or rely on skeletons to capture motion kinematics [Bianchi and Lanz(2025), Du et al.(2015)Du, Wang, and Wang, Duan et al.(2022)Duan, Zhao, Chen, Lin, and Dai]. Datasets like EgoExoLearn [Huang et al.(2024)Huang, Chen, Xu, Zhang, Yang, Pei, Zhang, Lu, Wang, Wang, and Qiao] and EgoExo4D [Grauman et al.(2024)Grauman, Westbury, Torresani, Kitani, Malik, Afouras, Ashutosh, Baiyya, Bansal, Boote, Byrne, Chavis, Chen, Cheng, Chu, Crane, Dasgupta, Dong, Escobar, Forigua, Gebreselasie, Haresh, Huang, Islam, Jain, Khirodkar, Kukreja, Liang, Liu, Majumder, Mao, Martin, Mavroudi, Nagarajan, Ragusa, Ramakrishnan, Seminara, Somayazulu, Song, Su, Xue, Zhang, Zhang, Castillo, Chen, Fu, Furuta, Gonzalez, Gupta, Hu, Huang, Huang, Khoo, Kumar, Kuo, Lakhavani, Liu, Luo, Luo, Meredith, Miller, Oguntola, Pan, Peng, Pramanick, Ramazanova, Ryan, Shan, Somasundaram, Song, Southerland, Tateno, Wang, Wang, Yagi, Yan, Yang, Yu, Zha, Zhao, Zhao, Zhu, Zhuo, Arbelaez, Bertasius, Damen, Engel, Farinella, Furnari, Ghanem, Hoffman, Jawahar, Newcombe, Park, Rehg, Sato, Savva, Shi, Shou, and Wray] emphasize the importance of multi-perspective inputs, but existing fusion strategies often rely on heuristic or late fusion, limiting generalization. In EgoExo4D [Grauman et al.(2024)Grauman, Westbury, Torresani, Kitani, Malik, Afouras, Ashutosh, Baiyya, Bansal, Boote, Byrne, Chavis, Chen, Cheng, Chu, Crane, Dasgupta, Dong, Escobar, Forigua, Gebreselasie, Haresh, Huang, Islam, Jain, Khirodkar, Kukreja, Liang, Liu, Majumder, Mao, Martin, Mavroudi, Nagarajan, Ragusa, Ramakrishnan, Seminara, Somayazulu, Song, Su, Xue, Zhang, Zhang, Castillo, Chen, Fu, Furuta, Gonzalez, Gupta, Hu, Huang, Huang, Khoo, Kumar, Kuo, Lakhavani, Liu, Luo, Luo, Meredith, Miller, Oguntola, Pan, Peng, Pramanick, Ramazanova, Ryan, Shan, Somasundaram, Song, Southerland, Tateno, Wang, Wang, Yagi, Yan, Yang, Yu, Zha, Zhao, Zhao, Zhu, Zhuo, Arbelaez, Bertasius, Damen, Engel, Farinella, Furnari, Ghanem, Hoffman, Jawahar, Newcombe, Park, Rehg, Sato, Savva, Shi, Shou, and Wray], baseline models are modality-specific—separate architectures are trained for first-person and third-person views, with predictions later combined via ensembling. These models also rely on heterogeneous pretraining datasets across modalities, making comparison and deployment under unified constraints more challenging.
Recent work on efficient adaptation, such as LoRA [Hu et al.(2021)Hu, Shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen], enables fine-tuning with fewer trainable parameters. Knowledge distillation also supports low-resource adaptation for vision tasks [Wang and Yoon(2022), Meneghetti et al.(2025)Meneghetti, Bianchi, Demo, and Rozza].
Our work advances the state-of-the-art by addressing key limitations in prior methods—such as reliance on modality-specific models, late fusion, and heavy pretraining. By introducing a unified architecture for multi-view skill classification, we simultaneously enhance classification accuracy and computational efficiency, enabling robust skill assessment under realistic resource constraints.
@@ 3 Proposed Methodology
We present SkillFormer, a novel architecture for unified multi-view proficiency estimation. Our model builds upon the TimeSformer [Bertasius et al.(2021)Bertasius, Wang, and Torresani] framework, incorporating parameter-efficient fine-tuning through LoRA [Hu et al.(2021)Hu, Shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen] and a specialized CrossViewFusion module for multi-view integration. Figure 1 provides an overview of our approach.
@@ 3.2 Video Feature Extraction
We use a TimeSformer backbone [Bertasius et al.(2021)Bertasius, Wang, and Torresani], pretrained on Kinetics-600 [Carreira and Zisserman(2017)], to extract spatio-temporal features by modeling spatial and temporal dependencies via divided attention. For efficient adaptation, we apply LoRA [Hu et al.(2021)Hu, Shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen] to key components—including attention projections, feed-forward layers, and temporal modules—allowing us to fine-tune only a small number of additional parameters.
@@
3.3 CrossViewFusion for Unified Multi-View Integration
We introduce CrossViewFusion, a lightweight yet expressive module for fusing information from multi-view video streams. It takes as input view-specific features from egocentric and exocentric sequences and applies a sequence of operations: normalization, cross-view attention, gated transformation, and adaptive self-calibration. These steps produce a unified embedding tailored for downstream expertise classification. An overview of the module is shown in Figure 2. Below, we describe each component of the CrossViewFusion module, aligned with its actual implementation.
View-Wise Normalization: To account for statistical disparities across viewpoints, we apply layer normalization independently to each view’s features. This operation ensures consistent feature scaling and centering, which is especially critical when merging modalities with inherently different distributions, such as first-person and third-person video streams.
Multi-Head Cross-View Attention: Rather than relying on simple concatenation or averaging, we employ a multi-head cross-attention mechanism, where each view can attend to and contextualize information from all other views. This enables the model to learn inter-view dependencies and adaptively prioritize view-specific cues. For instance, in cooking scenarios, egocentric hand movements may dominate attention, whereas in dance sequences, global body posture from exocentric views may become more salient.
Gated Feature Transformation The aggregated features are passed through a feed-forward network that includes a learnable gating mechanism. This consists of a linear layer followed by a sigmoid activation that produces gating weights for each hidden dimension. The gate acts as a dynamic filter, performing an element-wise modulation of the hidden representation to selectively amplify informative features while suppressing irrelevant ones. Such modulation is crucial for expertise assessment, where fine-grained differences in motion and control—not just coarse action labels—drive prediction quality.
Adaptive Feature Self-Calibration To further align the fused representations with the target classification space, we introduce an adaptive self-calibration block that learns per-feature affine parameters ($\mu_{\text{learn}},\sigma_{\text{learn}}$). This departs from traditional fixed-statistics normalization and allows the model to reshape feature distributions in a data-driven manner, improving its ability to resolve fine-grained skill variations across diverse activities.
Classification Head Finally, the calibrated features are fed into a classifier to predict discrete expertise levels.
@@ 4.1 Dataset
We evaluate on the Ego-Exo4D dataset [Grauman et al.(2024)Grauman, Westbury, Torresani, Kitani, Malik, Afouras, Ashutosh, Baiyya, Bansal, Boote, Byrne, Chavis, Chen, Cheng, Chu, Crane, Dasgupta, Dong, Escobar, Forigua, Gebreselasie, Haresh, Huang, Islam, Jain, Khirodkar, Kukreja, Liang, Liu, Majumder, Mao, Martin, Mavroudi, Nagarajan, Ragusa, Ramakrishnan, Seminara, Somayazulu, Song, Su, Xue, Zhang, Zhang, Castillo, Chen, Fu, Furuta, Gonzalez, Gupta, Hu, Huang, Huang, Khoo, Kumar, Kuo, Lakhavani, Liu, Luo, Luo, Meredith, Miller, Oguntola, Pan, Peng, Pramanick, Ramazanova, Ryan, Shan, Somasundaram, Song, Southerland, Tateno, Wang, Wang, Yagi, Yan, Yang, Yu, Zha, Zhao, Zhao, Zhu, Zhuo, Arbelaez, Bertasius, Damen, Engel, Farinella, Furnari, Ghanem, Hoffman, Jawahar, Newcombe, Park, Rehg, Sato, Savva, Shi, Shou, and Wray], a large-scale multimodal benchmark comprising over 1,200 hours of synchronized egocentric and exocentric video from 740 participants recorded across 123 real-world environments. Egocentric videos are first-person recordings captured from the demonstrator’s viewpoint, while exocentric videos are third-person views recorded from static cameras placed in the environment. The dataset spans eight domains—cooking, health, bike repair, music, basketball, bouldering, soccer, and dance—captured in naturalistic settings to reflect authentic task execution.
For the demonstrator proficiency estimation benchmark, only six domains are used: cooking, music, basketball, bouldering, soccer, and dance. The health and bike repair domains are excluded due to highly imbalanced or incomplete label distributions.
Each sample includes one egocentric video, recorded using Project Aria glasses [Engel et al.(2023)Engel, Somasundaram, Goesele, Sun, Gamino, Turner, Talattof, Yuan, Souti, Meredith, Peng, Sweeney, Wilson, Barnes, DeTone, Caruso, Valleroy, Ginjupalli, Frost, Miller, Mueggler, Oleinik, Zhang, Somasundaram, Solaira, Lanaras, Howard-Jenkins, Tang, Kim, Rivera, Luo, Dong, Straub, Bailey, Eckenhoff, Ma, Pesqueira, Schwesinger, Monge, Yang, Charron, Raina, Parkhi, Borschowa, Moulon, Gupta, Mur-Artal, Pennington, Kulkarni, Miglani, Gondi, Solanki, Diener, Cheng, Green, Saarinen, Patra, Mourikis, Whelan, Singh, Balntas, Baiyya, Dreewes, Pan, Lou, Zhao, Mansour, Zou, Lv, Wang, Yan, Ren, Nardi, and Newcombe], and up to four time-synchronized exocentric views captured via static GoPro cameras. Expert annotators assign one of four discrete proficiency levels—Novice, Early Expert, Intermediate Expert, or Late Expert—serving as ground truth for classification. The label distribution is notably imbalanced, skewed toward intermediate and late experts due to targeted recruitment of skilled participants.
We use all videos with available proficiency estimation labels, following the official benchmark’s training and validation splits. As in prior work [Braun et al.(2025)Braun, Armani, Meier, Moebus, and Holz], we train on the official training set, reserving 10% for validation, and evaluate on the full, held-out official validation set. This setup ensures direct and fair comparability with the accuracy reported in the dataset paper.
The dataset’s rich multi-view structure and diverse activity settings—ranging from kitchens and dance studios to climbing gyms and soccer fields—make it a strong testbed for real-world, multi-view skill assessment.
@@ 4.2 Implementation Details
All models were initialized from a TimeSformer backbone [Bertasius et al.(2021)Bertasius, Wang, and Torresani], pre-trained on Kinetics-600 [Carreira and Zisserman(2017)], and fine-tuned for 4 epochs using AdamW (weight decay = 0.01). We used a cosine annealing learning rate schedule with 1 epoch of linear warmup, a fixed batch size of 16, and 16 attention heads. Training was conducted on a single 80GB NVIDIA A100 GPU.
For each video, a fixed number of frames was sampled uniformly from a random temporal window using a frame sampling rate of $n$ (e.g., every $n$-th frame), ensuring temporal coverage within variable-length clips. The selected frames were resized such that the shortest edge is 224 pixels, center-cropped to $224\times 224$, rescaled to the $[0,1]$ range, and normalized with mean $[0.45,0.45,0.45]$ and standard deviation $[0.225,0.225,0.225]$.
@@
Table 1: Comparison with EgoExo4D proficiency estimation baselines (from [Grauman et al.(2024)Grauman, Westbury, Torresani, Kitani, Malik, Afouras, Ashutosh, Baiyya, Bansal, Boote, Byrne, Chavis, Chen, Cheng, Chu, Crane, Dasgupta, Dong, Escobar, Forigua, Gebreselasie, Haresh, Huang, Islam, Jain, Khirodkar, Kukreja, Liang, Liu, Majumder, Mao, Martin, Mavroudi, Nagarajan, Ragusa, Ramakrishnan, Seminara, Somayazulu, Song, Su, Xue, Zhang, Zhang, Castillo, Chen, Fu, Furuta, Gonzalez, Gupta, Hu, Huang, Huang, Khoo, Kumar, Kuo, Lakhavani, Liu, Luo, Luo, Meredith, Miller, Oguntola, Pan, Peng, Pramanick, Ramazanova, Ryan, Shan, Somasundaram, Song, Southerland, Tateno, Wang, Wang, Yagi, Yan, Yang, Yu, Zha, Zhao, Zhao, Zhu, Zhuo, Arbelaez, Bertasius, Damen, Engel, Farinella, Furnari, Ghanem, Hoffman, Jawahar, Newcombe, Park, Rehg, Sato, Savva, Shi, Shou, and Wray]). We report accuracy (%) for egocentric (Ego), exocentric (Exos), and combined views (Ego+Exos). SkillFormer outperforms the baselines in the Exos and Ego+Exos settings with significantly fewer trainable parameters and fewer epochs. Bold denotes the best accuracy; underlined values indicate second-best.
@@ 5 Results
We evaluate SkillFormer on the EgoExo4D [Grauman et al.(2024)Grauman, Westbury, Torresani, Kitani, Malik, Afouras, Ashutosh, Baiyya, Bansal, Boote, Byrne, Chavis, Chen, Cheng, Chu, Crane, Dasgupta, Dong, Escobar, Forigua, Gebreselasie, Haresh, Huang, Islam, Jain, Khirodkar, Kukreja, Liang, Liu, Majumder, Mao, Martin, Mavroudi, Nagarajan, Ragusa, Ramakrishnan, Seminara, Somayazulu, Song, Su, Xue, Zhang, Zhang, Castillo, Chen, Fu, Furuta, Gonzalez, Gupta, Hu, Huang, Huang, Khoo, Kumar, Kuo, Lakhavani, Liu, Luo, Luo, Meredith, Miller, Oguntola, Pan, Peng, Pramanick, Ramazanova, Ryan, Shan, Somasundaram, Song, Southerland, Tateno, Wang, Wang, Yagi, Yan, Yang, Yu, Zha, Zhao, Zhao, Zhu, Zhuo, Arbelaez, Bertasius, Damen, Engel, Farinella, Furnari, Ghanem, Hoffman, Jawahar, Newcombe, Park, Rehg, Sato, Savva, Shi, Shou, and Wray] proficiency estimation benchmark using classification accuracy as the primary evaluation metric, in line with the official proficiency demonstrator benchmark protocol. We consider three input configurations: egocentric (Ego, 1 view), exocentric (Exos, 4 views), and multi-view (Ego+Exos, 5 views), enabling us to isolate the contribution of each modality (Sections 5.1 and 5.2). Additionally, we analyze per-domain accuracy to evaluate generalization across diverse real-world tasks (Section 5.3). SkillFormer achieves consistent improvements in both accuracy and efficiency, outperforming multi-view baselines with fewer parameters and significantly lower training cost.

### v3

# Further author information: (Send correspondence to E.B.)

@@ Abstract
Assessing human skill levels in complex activities is a challenging problem with applications in sports, rehabilitation, and training. In this work, we present SkillFormer, a parameter-efficient architecture for unified multi-view proficiency estimation from egocentric and exocentric videos. Building on the TimeSformer backbone, SkillFormer introduces a CrossViewFusion module that fuses view-specific features using multi-head cross-attention, learnable gating, and adaptive self-calibration. We leverage Low-Rank Adaptation to fine-tune only a small subset of parameters, significantly reducing training costs. In fact, when evaluated on the EgoExo4D dataset, SkillFormer achieves state-of-the-art accuracy in multi-view settings while demonstrating remarkable computational efficiency, using 4.5x fewer trainable parameters and requiring 3.75x fewer training epochs than prior baselines. It excels in multiple structured tasks, confirming the value of multi-view integration for fine-grained skill assessment.
keywords:
Video Understanding, Proficiency Estimation, Action Quality Assessment
@@ 1 Introduction
To address this, we propose SkillFormer, a lightweight and efficient architecture for multi-view skill assessment. SkillFormer processes synchronized egocentric and exocentric video streams using a TimeSformer backbone, enhanced by a novel CrossViewFusion module that leverages cross-attention and learnable gating. To enable efficient fine-tuning, it incorporates Low-Rank Adaptation (LoRA). This design achieves competitive overall accuracy while improving computational efficiency through reduced trainable parameters and faster convergence.
@@ 2 Background and Related Works
Action Quality Assessment (AQA) focuses on evaluating human action quality with applications in sports and training [1]. Traditional approaches used handcrafted features or 3D CNNs [2, 3, 4], while recent methods adopt transformer architectures like TimeSformer and Video Swin for long-range dependencies [5, 6]. Specialized designs include GDLT’s grade-aware features [7] and CoRe’s contrastive regression [8].
Multi-modal AQA improves robustness by integrating complementary signals. Some approaches combine RGB and depth for industrial tasks [9], while others rely on skeletons to capture motion kinematics [10, 11]. Datasets like EgoExo4D [12] provides multi-perspective data but baselines use separate architectures with heterogeneous pretraining, limiting unified deployment
Parameter-Efficient Adaptation through LoRA [13] enables fine-tuning with reduced parameters, addressing computational constraints in video understanding tasks. Knowledge distillation also supports low-resource adaptation for vision tasks [14].
Our work addresses these limitations through unified multi-view architecture with efficient fusion and consistent pretraining, simultaneously improving accuracy and computational efficiency.
@@ 3 Proposed Methodology
We present SkillFormer, a novel architecture for unified multi-view proficiency estimation. Our model builds upon the TimeSformer [5] framework, incorporating parameter-efficient fine-tuning through LoRA [13] and a specialized CrossViewFusion module for multi-view integration. Figure 1 provides an overview of our approach.
@@ 3.2 Video Feature Extraction
We use a TimeSformer backbone [5], pretrained on Kinetics-600 [3], to extract spatio-temporal features by modeling spatial and temporal dependencies via divided attention. For efficient adaptation, we apply LoRA [13] to key components—attention projections (qkv), attention output dense layers, temporal attention components, and feed-forward layers—allowing us to fine-tune only a small number of parameters. During inference, LoRA adapters are merged with the base model weights, eliminating computational overhead while maintaining full model capacity.
@@
Figure 2: Detailed architecture of the CrossViewFusion module. Input features $(B,V,d)$ undergo: (1) Layer normalization per view, (2) Multi-head cross-attention enabling each view to attend to all others, (3) View aggregation via mean pooling, (4) Feed-forward transformation with GELU activation, (5) Learnable gating mechanism $\mathbf{g}=\sigma(\text{Linear}(\mathbf{h}))$ for selective feature modulation, (6) Final projection, and (7) Adaptive self-calibration using learnable statistics to align with classification space.
3.3 CrossViewFusion for Unified Multi-View Integration
We introduce CrossViewFusion, a lightweight module for fusing multi-view video streams through normalization, cross-view attention, gated transformation, and adaptive self-calibration (Figure 2).
Given view-specific features $\mathbf{X}\in\mathbb{R}^{B\times V\times d}$ where $B$ is batch size, $V$ is number of views, and $d=768$, the module applies:
View-wise Normalization and Multi-Head Cross-View Attention: To account for statistical disparities across viewpoints, we first apply layer normalization independently to each view’s features. We then employ multi-head attention where each view attends to all views (including itself), enabling the model to learn inter-view dependencies and adaptively weight view-specific cues:
$\mathbf{X}_{norm}$ $=\text{LayerNorm}(\mathbf{X})$ (1)
@@ 3.3 CrossViewFusion for Unified Multi-View Integration
View Aggregation and Feature Transformation: We aggregate the attention-weighted features through averaging and pass them through a feed-forward network with GELU activation, layer normalization, and dropout for regularization:
$\mathbf{h}$ $=\frac{1}{V}\sum_{v=1}^{V}\mathbf{X}_{attn}[:,v,:]$ (3)
$\mathbf{h}_{proj1}$ $=\text{Linear}_{1}(\mathbf{h})$ (4)
$\mathbf{h}_{hidden}$ $=\text{Dropout}(\text{LayerNorm}(\text{GELU}(\mathbf{h}_{proj1})))$ (5)
Learnable Gating: The aggregated features are modulated through a learnable gating mechanism that produces element-wise weights via a sigmoid activation. This gate acts as a dynamic filter to selectively amplify informative features while suppressing irrelevant ones:
$\mathbf{g}$ $=\sigma(\text{Linear}_{gate}(\mathbf{h}_{hidden}))$ (6)
$\mathbf{h}_{gated}$ $=\mathbf{g}\odot\mathbf{h}_{hidden}$ (7)
Final Projection and Adaptive Self-Calibration: To align the fused representations with the target classification space, we apply a final linear projection followed by adaptive self-calibration. Unlike traditional fixed-statistics normalization, we learn per-feature affine parameters to reshape feature distributions in a data-driven manner:
$\mathbf{h}_{proj2}$ $=\text{Linear}_{2}(\mathbf{h}_{gated})$ (8)
$\mathbf{h}_{norm}$ $=\text{LayerNorm}(\mathbf{h}_{proj2})$ (9)
$\mathbf{h}_{centered}$ $=\mathbf{h}_{norm}-\text{mean}(\mathbf{h}_{norm},\text{dim}=-1)$ (10)
$\mathbf{h}_{scaled}$ $=\frac{\mathbf{h}_{centered}}{\text{std}(\mathbf{h}_{norm},\text{dim}=-1)+\epsilon}$ (11)
$\mathbf{h}_{final}$ $=\mathbf{h}_{scaled}\cdot\sigma_{learn}+\mu_{learn}$ (12)
where $\mu_{learn},\sigma_{learn}$ are learnable statistics that allow the model to adaptively reshape feature distributions for improved skill assessment across diverse activities.
@@ 4.1 Dataset
We evaluate on the Ego-Exo4D dataset [12], a large-scale benchmark with over 1,200 hours of synchronized egocentric (first-person) and exocentric (third-person) video from 740 participants across 123 real-world environments. Each sample includes one egocentric video (Project Aria glasses) and up to four time-synchronized exocentric views (static GoPro cameras).
For proficiency estimation, we use six domains (cooking, music, basketball, bouldering, soccer, dance) with four discrete levels: Novice, Early Expert, Intermediate Expert, and Late Expert. Health and bike repair domains are excluded due to imbalanced label distributions. The dataset is skewed toward intermediate and late experts due to targeted recruitment of skilled participants.
We use all videos with available proficiency estimation labels, following the official benchmark’s training and validation splits. As in prior work [15], we train on the official training set, reserving 10% for validation, and evaluate on the full, held-out official validation set. This setup ensures direct and fair comparability with the accuracy reported in the dataset paper.
@@ 4.2 Implementation Details
Models use a TimeSformer backbone [5] pretrained on Kinetics-600 [3], fine-tuned for 4 epochs with AdamW (weight decay=0.01), cosine annealing schedule, batch size 16, on single A100 GPU.
Videos are center-cropped to $\times 224$ 224, rescaled to the [0,1][0,1][ 0 , 1 ] range, and normalized with mean 0.450.450.45 and standard deviation 0.2250.2250.225. Frames are sampled uniformly with temporal interpolation of positional embeddings for variable lengths.
@@
Table 1: Comparison with EgoExo4D proficiency estimation baselines [12] and with EgoPulseFormer [15], a concurrent approach. We report accuracy (%) for egocentric (Ego), exocentric (Exos), and combined views (Ego+Exos). SkillFormer outperforms the baselines in the Exos and Ego+Exos settings with significantly fewer trainable parameters and fewer epochs. Bold denotes the best accuracy; underlined values indicate second-best.
@@ 5 Results
We evaluate SkillFormer on the EgoExo4D [12] proficiency estimation benchmark using classification accuracy as the primary evaluation metric, in line with the official proficiency demonstrator benchmark protocol. We consider three input configurations: egocentric (Ego, 1 view), exocentric (Exos, 4 views), and multi-view (Ego+Exos, 5 views), enabling us to isolate the contribution of each modality (Sections 5.1 and 5.2). Additionally, we analyze per-domain accuracy to evaluate generalization across diverse real-world tasks (Section 5.3). SkillFormer achieves consistent improvements in both accuracy and efficiency, outperforming multi-view baselines with fewer trainable parameters and significantly lower training cost.